OpenStudy (anonymous):

8/9 + 1/3 / 6/7 simplest form

OpenStudy (lizz123):

add 8/9 + 1/3 = 11/9 = 1 2/9 1 2/9 /67

OpenStudy (lizz123):

Once I divided I got the answer of 1 23/54

OpenStudy (lizz123):

 1 23/54 is you answer